# EP 250108a/SN 2025kg: Observations of the most nearby Broad-Line Type Ic Supernova Following an Einstein Probe Fast X-ray Transient

### v1

@@ Abstract
With a small sample of fast X-ray transients (FXTs) with multi-wavelength counterparts discovered to date, the progenitors of FXTs and their connections to -ray bursts (GRBs) and supernovae (SNe) remain ambiguous. Here, we present photometric and spectroscopic observations of SN 2025kg, the supernova counterpart to the FXT EP 250108a. At , this is the closest known SN discovered following an Einstein Probe (EP) FXT. We show that SN 2025kg’s optical spectra reveal the hallmark features of a broad-lined Type Ic SN. Its light curve evolution and expansion velocities are comparable to those of GRB-SNe, including SN 1998bw, and two past FXT SNe. We present JWST/NIRSpec spectroscopy taken around SN 2025kg’s maximum light, and find weak absorption due to He I m and a broad, unidentified emission feature at 4–4.5 m. Further, we observe clear evidence for broadened H in optical spectra at 42.5 days that is not detected at other epochs, indicating interaction with H-rich material. From its light curve, we derive a 56Ni mass of 0.2 - 0.6 Together with our companion paper (Eyles-Ferris et al., 2025), our broadband data of EP 250108a/SN 2025kg are consistent with a trapped or low energy ( ergs) jet-driven explosion from a collapsar with a zero-age main sequence mass of 15–30 . Finally, we show that the sample of EP FXT SNe support past rate estimates that low-luminosity jets seen through FXTs are more common than successful (GRB) jets, and that similar FXT-like signatures are likely present in at least a few percent of the brightest Ic-BL SNe.
@@ 4.1 One-zone Model
Our model provides a strong fit to the data. We infer of ejecta with a 56Ni mass of and an initial ejecta velocity of km s-1. Our estimate of ejecta mass and energy is consistent with median values inferred from Type Ic-BL SNe (Taddia et al., 2019) using a similar model. However, our estimate of the 56Ni is marginally inconsistent with the Type Ic-BL SNe medial values (at ), which could be due to differences in assumptions about -ray opacities or hinting towards a limitation with the one-zone model.

### v2

@@ Abstract
With a small sample of fast X-ray transients (FXTs) with multi-wavelength counterparts discovered to date, their progenitors and connections to -ray bursts (GRBs) and supernovae (SNe) remain ambiguous. Here, we present photometric and spectroscopic observations of SN 2025kg, the SN counterpart to the FXT EP 250108a. At , this is the closest known SN discovered following an Einstein Probe (EP) FXT. We show that SN 2025kg’s optical spectra reveal the hallmark features of a broad-lined Type Ic SN. Its light curve evolution and expansion velocities are comparable to those of GRB-SNe, including SN 1998bw, and two past FXT SNe. We present JWST/NIRSpec spectroscopy taken around SN 2025kg’s maximum light, and find weak absorption due to He I m and a broad, unidentified emission feature at 4–4.5 m. Further, we observe broadened H in optical data at 42.5 days that is not detected at other epochs, indicating interaction with H-rich material. From its light curve, we derive a 56Ni mass of 0.2 - 0.6 . Together with our companion paper (Eyles-Ferris et al., 2025), our broadband data are consistent with a trapped or low energy ( ergs) jet-driven explosion from a collapsar with a zero-age main sequence mass of 15–30 . Finally, we show that the sample of EP FXT SNe support past estimates that low-luminosity jets seen through FXTs are more common than successful (GRB) jets, and that similar FXT-like signatures are likely present in at least a few percent of the brightest Ic-BL SNe.
@@ 4.1 One-zone Model
Our model provides a strong fit to the data. We infer of ejecta with a 56Ni mass of and an initial ejecta velocity of km s-1. Our estimate of ejecta mass and energy is consistent with median values inferred from Type Ic-BL SNe (Taddia et al., 2019; Srinivasaragavan et al., 2024) using a similar model. However, our estimate of the 56Ni is marginally inconsistent with the Type Ic-BL SNe medial values of Taddia et al. (2019) (at ), which could be due to differences in assumptions about -ray opacities or hinting towards a limitation with the one-zone model.